Lawrence Corporation is considering these three projects whose returns are normally distributed:

Project Investment Exp. Ret Std. Dev. Corr. Coeff.
A $20,000 12% 20% A, B = 0.4
B $30,000 14% 20% A, C = 0.5
C $50,000 16% 30% B, C = 0.6

Find the probability that the return on the portfolio is more than 20%
Answer #1
a b=a/100000
Investment Weight
Project A $20,000               0.20 wa
Project B $30,000               0.30 wb
Project C $50,000               0.50 wc
$100,000
Ra=Return of Project A 12%
Rb=Return of Project B 14%
Rc=Return of Project C 16%
Portfolio Return =wa*Ra+wb*Rb+wc*Rc
Portfolio Return =0.2*12+0.3*14+0.5*16= 14.60%
Sa=Standard Deviation of Project A 20%
Sb=Standard Deviation of Project B 20%
Sc=Standard Deviation of Project C 30%
Corr(a,b)=Correlation of Project A and B           0.40
Corr(a,c)=Correlation of Project A and C           0.50
Corr(b,c)=Correlation of Project B and C           0.60
Covariance (a,b)=Correlation(a,b)*Sa*Sb      160.00 %% (0.4*20*20)
Covariance (a,c)=Correlation(a,c)*Sa*Sc      300.00 %% (0.5*20*30)
Covariance (b,c)=Correlation(b,c)*Sb*Sc      360.00 %% (0.6*20*30)
Portfolio Variance =(wa^2)*(Sa^2)+(wb^2)*(Sb^2)+(wc^2)*(Sc^2)+2wa*wb*Cov(a,b)+2wa*wc*Cov(a,c)+2wb*wc*Cov(b,c)
Portfolio Variance 464.2 %% (0.2^2)*(20^2)+(0.3^2)*(20^2)+(0.5^2)*(30^2)+2*0.2*0.3*160+2*0.2*0.5*300+2*0.3*0.5*360
Portfolio Standard Deviation =SQUARE ROOT (Variance)
Portfolio Standard Deviation 21.5453 % SQRT(464.2)
Portfolio Mean Return 14.60%
Probability of return of the portfolio more than 20%
Portfolio Variance =(wa^2)*(Sa^2)+(wb^2)*(Sb^2)+(wc2)*(Sc^2)+2wa wb*Cov(a,b)+2wa wc Covía,c)+2wb*wc*Cov(b,c) 464.2 %% (0.242)